# Beyond the Linear Separability Ceiling

### v1

@@ Abstract
Most state-of-the-art Visual-Language Models (VLMs) are seemingly limited by the linear separabilty of their visual embeddings on abstract reasoning tasks. This work investigates this “linear reasoning bottleneck” by introducing the Linear Separability Ceiling (LSC), the performance of a simple linear classifier on a VLM’s visual embeddings. We find this bottleneck is widespread and stems not from poor perception, but from failures in the language model’s reasoning pathways. We demonstrate this is a solvable alignment issue. The required intervention, however, is task-dependent: activating existing pathways suffices for semantic concepts, while complex relational reasoning requires adapting core model weights. Using postfix tuning as a methodological control, we find strong evidence for powerful, dormant reasoning pathways within VLMs. However, for complex relational tasks requiring deeper adaptation, explicitly improving representation quality causes the model to fail on new prompt formats despite its embeddings remaining well separated. Ultimately, this work provides a new lens for VLM analysis, showing that robust reasoning is a matter of targeted alignment, not simply improved representation learning.
@@ 1 Introduction
A critical challenge in advancing state-of-the-art Visual-Language Models (VLMs) (Radford et al., 2021) lies in understanding the root of their frequent failures on abstract tasks. Is the bottleneck poor visual perception or flawed reasoning? This unresolved question creates a persistent gap between machine and human cognition, particularly on visual puzzles (Wüst et al., 2024) and tasks that VLMs can solve in text but not in visual formats (Park et al., 2025). While previous studies have explored this perception-reasoning interface, they have not definitively pinpointed the primary cause of failure (Vaishnav & Tammet, 2025; Lin et al., 2025).
To address this ambiguity, this paper introduces a diagnostic framework centered on the Linear Separability Ceiling (LSC): a measure of the maximum performance a linear classifier can achieve on a VLM’s raw visual embeddings. The LSC provides a clear baseline for the quality of the initial visual representations, establishing a benchmark that the model’s end-to-end reasoning must surpass to demonstrate added value.
Our focus on computational non-linearity is complemented by recent work on representational geometry by Engels et al. (2025) that challenges the Linear Representation Hypothesis (LRH). They find that concepts like days of the week and months of the year form irreducible, circular features in activation space. This representational non-linearity implies a need for the computational non-linearity we investigate; as they show, models use these exact circular features to perform tasks involving modular arithmetic. Their work thus offers a concrete, mechanistic example of the structured features our hypothesized reasoning pathways may operate on.
Applying this framework reveals our central finding: a widespread "linear reasoning bottleneck," where the generative performance of most leading VLMs is statistically equivalent or lower than the LSC on their own visual embeddings. The few models that surpass this ceiling do so via two distinct pathways: either by refining embeddings into a more linearly separable format or by applying effective non-linear reasoning.
The discovery — that reasoning, not perception, is often the main bottleneck — motivates our investigation into unlocking this latent potential. We show the reasoning bottleneck is a solvable alignment issue. Through targeted parameter-efficient fine-tuning (PEFT) (Hu et al., 2022), we enable models to significantly surpass their initial LSC. The success of postfix tuning (Li & Liang, 2021; Lester et al., 2021), a methodological control that cannot alter visual representations, provides strong evidence that VLMs possess powerful, dormant reasoning pathways.
@@ 2 Related work
Abstract visual reasoning in VLMs. The evaluation of Vision-Language Models (VLMs) has evolved through a series of increasingly complex benchmarks. Foundational benchmarks focused on single-image understanding, including image captioning (Lin et al., 2014), Optical Character Recognition (OCR) (Mishra et al., 2019), and Visual Question Answering (VQA) (Hudson & Manning, 2019; Goyal et al., 2019). To assess multi-image understanding with language Meng et al. (2025) developed the Multimodal Multi-Image Understanding (MMIU) dataset. To evaluate core visual understanding beyond language, researchers use purely visual reasoning benchmarks that require identifying abstract concepts and relations. This category includes Bongard tasks (Bongard, 1970), which we focus on. But also Raven’s Progressive Matrices (RPMs) (Zhang et al., 2019), Visual Progressive Matrices (V-PROM) (Teney et al., 2020), Compositional Visual Reasoning (CVR) (Zerroug et al., 2022), and the Abstract Reasoning Corpus (ARC) (Chollet, 2019; Unsal & Akkus, 2025).
@@
VLM architectures and modality fusion. A key architectural differentiator in VLMs is the strategy for fusing modalities (Gadzicki et al., 2020; Shukor et al., 2025). While early-fusion models like Chameleon (Team, 2024) create a joint representation from raw inputs, the now-dominant late-fusion approach first processes images with a dedicated vision encoder. In its most common form, exemplified by LLaVA (Liu et al., 2023), the resulting visual embeddings are mapped into the language model’s space via a simple projection layer. A more deeply integrated variant is cross-attention fusion, pioneered by Flamingo (Alayrac et al., 2022) and BLIP (Li et al., 2022), which insert cross-attention layers within the LLM, allowing textual tokens to dynamically attend to visual features throughout the model for a more iterative fusion (Lin et al., 2021). Our framework is designed for late-fusion models where initial visual representation can be isolated.
Mechanistic interpretability aims to reveal the internal workings of neural networks, explaining how specific outputs are generated. Common approaches include probing, which uses simple auxiliary classifiers on intermediate outputs to identify encoded information; logit lens analysis, which examines layer-by-layer prediction changes; and analysis of sparse autoencoders or attention mechanisms. (Liu et al., 2025a; Golovanevsky et al., 2025) Our framework applies the probing paradigm to assess how concept separability evolves across the model’s internal representations.
@@
Linear representation hypothesis (LRH) posits that concepts within neural networks are represented by linear structures in activation space. This is commonly observed as compositional linearity, where the embeddings of composite concepts can be decomposed into the sum of their constituent parts (Trager et al., 2023). This principle appears to be a fundamental aspect of transformer decoders, where even the transformations between sequential layers exhibit near-perfect linearity (Razzhigaev et al., 2024). Theoretical work supports these empirical findings, suggesting that the next-token prediction objective inherently biases models towards learning representations that are linear transformations of latent concepts (Liu et al., 2025b).
@@
Representation alignment and the modality gap research has identified critical bottlenecks that can occur either in the initial visual encoding or in how the language model component subsequently extracts and utilizes these visual features for complex reasoning tasks (Chia et al., 2024; Zhang et al., 2024). Compounding these challenges is the modality gap, a widely observed phenomenon where the learned representations for different modalities exhibit distinct structures or occupy separate regions within the joint embedding space, even when they are semantically aligned (Yaras et al., 2022). This systematic separation, particularly noted in contrastively trained models, can hinder cross-modal alignment and the seamless integration of information (Qiu et al., 2024), potentially impacting the discriminative power of these representations. Our framework allows us to quantitatively assess the impact of such issues by establishing the linear separability ceiling as baseline. This concept of a representational performance ceiling is not new; for instance, prior work has identified a “contrastive learning ceiling” in semantic textual similarity (Zhang & Li, 2024).
@@
Training and adapting vision encoders in VLMs are typically pre-trained with methods such as contrastive objectives (Chen et al., 2020), exemplified by CLIP (Radford et al., 2021), or combined contrastive and generative objectives during pretraining, characteristic of CoCa (Yu et al., 2022). Building on this, some approaches experiment with contrastive losses to augment NTP to align VLMs (Ouali et al., 2025; Wu et al., 2025; Ak et al., 2024). To efficiently implement such strategies, PEFT methods like Low-Rank Adaptation (LoRA) (Hu et al., 2022) and prompt tuning (Li & Liang, 2021; Zhou et al., 2022) are widely adopted for efficiently adapting VLMs (Lester et al., 2021; Jahan et al., 2025). We investigate whether strategies that explicitly co-optimize for enhanced representational discriminability alongside generative accuracy enable VLMs to better leverage their internal visual representations.
@@ 3 Experimental setup
Figure 1: An example of the bongard HOI task. Model must infer a common rule, in this case, "a person is performing a jump on a motorcycle," from the positive examples and determine if the query image follows this rule, which the negative examples do not.
Datasets. We conduct experiments on two such datasets: Bongard OpenWorld (Wu et al., 2024) and Bongard Human-Object-Interaction (HOI) (Jiang et al., 2022). The OpenWorld dataset was divided into train (500), validation (100), and test (500) sample splits, ensuring distinct semantic components across them. For HOI, we utilized its original splits, where 4000 balanced samples across concepts were chosen from train split. Its validation ($4\cdot 100$) and test sets ($4\cdot 200$) balanced samples, which are categorized by concept novelty: seen/unseen object/action. Each sample consisted of 6 positive images, 6 negative images and 1 query image.
Models. Baseline performance was established on multiple SoTA VLMs: Phi 3.5 vision 4.2B (Microsoft, 2024), Pixtral 12B (MistralAI, 2024), Gemma3 4B and 27B (Google, 2025), InternVL3 14B (Chen et al., 2024), and Qwen 2.5 VL 7B and 72B (Alibaba, 2025). Subsequently, we applied PEFT to Gemma3 4B, Phi and Pixtral.
Evaluation. For all fine-tuning experiments, we saved model checkpoints every epoch and selected the one with the best performance on the validation set. Baselines and best-performing checkpoints were evaluated on their respective test splits. Models demonstrate consistent performance across all conceptual novelty splits (see Appendix I), so we report the average for brevity.
@@ 4 A framework for decomposing VLM reasoning
To diagnose whether VLM failures on abstract reasoning tasks stem from flawed perception or reasoning, we introduce a diagnostic framework centered on a non-parametric linear probe. This framework assesses the discriminability of a model’s visual embeddings, independent of its generative process, allowing us to disentangle the contributions of representation quality from non-linear reasoning. First, we benchmark performance across various prompt structures, including interleaved vs. labeled image presentation and direct vs. Chain-of-Thought (CoT) prompting (see Appendix A).
@@ 4.1 Methodology
We first extract multi-token image embeddings at two key stages: the vision stage, which captures initial representations from the vision encoder before contextualization, and the final stage, which captures contextualized representations from the LLM’s final hidden state. We aggregate these sequences into single vectors, $\vec{v}_{i}$, using mean pooling and L2 normalization. While other methods like max pooling were explored, mean pooling is an effective, computationally efficient approximation of slerp (Reimers & Gurevych, 2019), which preserves the angular properties measured by cosine similarity (Shoemake, 1985; Karras et al., 2019). To quantify the discriminative quality of these embeddings, we probe them using a nearest-centroid classifier. We selected this nearest-centroid strategy as it proved more performant than a nearest-neighbor approach because it mitigates the influence of outlier examples as demonstrated in Appendix B (Hastie et al., 2009). We then compute prototype vectors for the positive ($P$) and negative ($N$) example sets by averaging their respective image vectors (e.g., $\vec{c}_{P}=\text{mean}(\vec{v}_{p_{1}},\dots,\vec{v}_{p_{k}})$). A query image ($Q$) is then classified based on its cosine similarity to these centroids (i.e., comparing $\cos(\vec{v}_{Q},\vec{c}_{P})$ to $\cos(\vec{v}_{Q},\vec{c}_{N})$).
4.2 Decomposition of performance
@@
Our framework dissects VLM performance by statistically comparing three key accuracies (Figure 2): the end-to-end generative accuracy ($acc_{gen}$), the linear probe accuracy on initial vision embeddings ($acc_{vision}$), and on final contextualized embeddings ($acc_{final}$). Grounded in the linear representation hypothesis, which posits that concepts are encoded as linear structures within a model’s activation space, we define $acc_{vision}$ as the linear separability ceiling (LSC). The LSC quantifies the baseline discriminative power of the visual representations, a primary goal of contrastive pre-training. While we recognize this linear assumption is a simplification that may not capture all complex concept geometries, it establishes the benchmark a model must surpass to demonstrate effective non-linear reasoning. We also note that accuracy can also be misleading under class imbalance, as the embeddings’ discriminative power often exhibits class-dependent variability (Appendix K).
Figure 2: Flowchart illustrating the derivation of the $acc_{vision}$ (LSC), $acc_{final}$, and $acc_{gen}$ metrics.
We classify model performance based on a statistical comparison (non-overlapping 95% confidence intervals) between $acc_{gen}$ and the LSC. Generative performance is surpassed if $acc_{gen}$ is superior to the LSC, linear reasoning bottleneck if $acc_{gen}$ is statistically indistinguishable or lower than LSC. For models that surpass the LSC, we diagnose the mechanism by comparing $acc_{gen}$ to $acc_{final}$: if they are indistinguishable, success is due to non-linear representation refinement; if $acc_{gen}$ is superior, it points to post-representation non-linear reasoning. A Chi-squared ($\chi^{2}$) test assesses the statistical dependence between the generative model and our linear probes.
5 Linear reasoning bottleneck

### v3

@@ Abstract
A challenge in advancing Visual-Language Models (VLMs) is determining whether their failures on abstract reasoning tasks, such as Bongard problems, stem from flawed perception or faulty top-down reasoning. To disentangle these factors, we introduce a diagnostic framework centered on the Linear Separability Ceiling (LSC), the performance achievable by a linear classifier on a VLM’s raw visual embeddings. Applying this framework to state-of-the-art VLMs, we uncover a pervasive “alignment gap”, where most models fail to generatively outperform the linear separability of their representations. We find that the few models surpassing this ceiling do so via two mechanisms: by further refining visual representations into a more linearly separable format or by executing non-linear decision logic. We demonstrate that this bottleneck is not a fundamental limitation but a solvable visual alignment issue. Our method augments standard next-token prediction with a contrastive objective to restructure the visual manifold into a more one-dimensionally linear geometry, improving image-to-image comparison and enabling models to significantly surpass the LSC on abstract binary classification tasks.
@@ 1 Introduction
A challenge for state-of-the-art Visual-Language Models (VLMs) (radford2021learning) is understanding the root of their frequent failures on abstract tasks. Is the bottleneck flawed bottom-up visual perception, which builds representations from sensory input (marr1982vision), or flawed top-down reasoning that interprets that input using prior knowledge (gregory1970intelligent)? This question highlights a persistent gap between machine and human cognition, particularly on visual puzzles (wst2024bongard) and tasks VLMs solve in text but not in visual formats (park2025generalizing). While previous studies have explored this perception-reasoning interface within modern VLMs for abstract tasks like Bongard problems, the primary cause of failure has been difficult to isolate, lacking an empirical method to quantify this gap (vaishnav2025a; limitations2025miko).
To address this ambiguity, this paper introduces a diagnostic framework centered on the Linear Separability Ceiling (LSC), a measure of the performance a linear classifier can achieve on a VLM’s raw visual embeddings. The LSC provides a baseline for the quality of the initial visual representations, establishing a benchmark that the model’s end-to-end (non-linear) reasoning must surpass to demonstrate added value.
Our focus on computational non-linearity is complemented by recent work on representational geometry by engels2025not that challenges the Linear Representation Hypothesis (LRH). They find that concepts like days of the week and months of the year form irreducible, circular features in activation space. This representational non-linearity implies a need for the computational non-linearity we investigate; as they show, models use these exact circular features to perform tasks involving modular arithmetic. Their work thus offers a concrete, mechanistic example of the structured features our hypothesized (non-linear) reasoning pathways may operate on.
Applying this framework reveals our central finding: an “alignment gap,” a concept that extends known representational issues like the “modality gap” (yaras2022mind) to the interface between perception and reasoning. We find that a model’s reasoning pathways are often not aligned with its own high-quality visual representations. This causes the generative performance of most leading VLMs to be statistically no better than the LSC on their own visual embeddings.
The models that surpass this ceiling do so via two strategies: by further refining representations into a more linearly separable format, effectively extending the bottom-up perception process, or by executing a non-linear decision logic that functions as a form of top-down reasoning. Strong evidence for this second pathway comes from the success of postfix tuning (li2021prefixtuning; lester2021the), a method which by design cannot alter the initial representations.
@@ 2 Related work
Abstract visual reasoning in VLMs. The evaluation of VLMs has evolved from foundational benchmarks for single-image understanding, such as captioning (lin2014microsoft) and VQA (goyal2017making), to more rigorous tests of reasoning. The complexity of these evaluations has increased along two axes. First, at the vision-language interface, benchmarks now demand deeper reasoning. This includes VQA variants that test for compositional structure and spatial skills like GQA (hudson2019gqa), or probe commonsense understanding as in VCR (zellers2019vcr). Complementing these, datasets like Winoground (thrush2022winoground) specifically isolate visio-linguistic compositional abilities. Second, in the purely visual domain, a distinct set of benchmarks evaluates abstract reasoning by removing the linguistic component entirely. This category includes Raven’s Progressive Matrices (RPMs) (zhang2019raven), the Abstract Reasoning Corpus (ARC) (chollet2019arc), and Bongard tasks (bongard1970pattern), which are the focus of this work.
@@
VLM architectures and modality fusion. A key architectural differentiator in VLMs is the strategy for fusing modalities (gad2020early; mustafa2025scaling). While early-fusion models like Chameleon (chameleonteam2024chameleon) create a joint representation from raw inputs, the now-dominant late-fusion approach first processes images with a dedicated vision encoder. In its most common form, exemplified by LLaVA (liu2023visual), the resulting visual embeddings are mapped into the language model’s space via a simple projection layer. A more deeply integrated variant is cross-attention fusion, pioneered by Flamingo (alayrac2022flamingo) and BLIP (li2022blip), which insert cross-attention layers within the LLM (lin2021cat). Our framework is designed for late-fusion models where initial visual representation can be isolated.
@@
Linear Representation Hypothesis (LRH) posits that concepts within neural networks are represented by linear structures in activation space. This is commonly observed as compositional linearity, where the embeddings of composite concepts can be decomposed into the sum of their constituent parts (trager2023linear), and appears fundamental to transformer decoders, whose sequential layer transformations are often nearly linear (razzhigaev2024your). Theoretical work supports these empirical findings, suggesting that the next-token prediction objective inherently biases models towards learning representations that are linear transformations of latent concepts (liu2025i). This hypothesis suggests that the learned representations should be linearly separable — that is, distinguishable by a simple linear classifier. This property is critical for generalization, enabling linear transferability where a classifier trained on a source domain remains effective on a new, related target domain (haochen2022separability). The emergence of such a well-ordered geometric space is not a coincidence but a predictable consequence of high-dimensional geometry. Stochastic separation theorems establish that in high dimensions, any given point in a random set can be separated from the others by a hyperplane with high probability, even if the number of points in the set grows exponentially with the dimension (sidorov2020).
@@
Representation alignment and the modality gap. Research into representation alignment has identified critical bottlenecks in VLMs, which can occur either in the initial visual encoding or in the subsequent reasoning process (chia2024puzzlevqa; zhang2024can). A key example of such an alignment challenge is the modality gap: a widely observed phenomenon where the learned representations for different modalities exhibit distinct structures or occupy separate regions within the joint embedding space, even when they are semantically aligned (yaras2022mind). This systematic separation, particularly noted in contrastively trained models, can hinder cross-modal alignment and the seamless integration of information (qiu2024mining), potentially impacting the discriminative power of these representations. Our framework allows us to quantitatively assess the impact of such issues by establishing the linear separability ceiling as baseline. This concept of a representational performance ceiling is not new; for instance, prior work has identified a “contrastive learning ceiling” in semantic textual similarity (zhang2024pcctuning).
@@
Training and adapting. To bridge the modality gap and enhance reasoning, various adaptation strategies have been developed. While some work proposes novel architectural modifications (bigverdi2024perception; kolner2025mind) or reinforcement learning techniques (li2025selfrewarding; huang20253d), our work focuses on training objectives. Vision encoders are typically pre-trained with objectives like contrastive learning, as seen in CLIP (radford2021learning), or combined contrastive-generative objectives, as in CoCa (yu2022coca). Building on this, some methods augment the standard next-token prediction objective with an auxiliary contrastive loss to explicitly align representations (ouali2025vladva; wu2025symmetrical; kenanemirak2024aligning). While these approaches typically tackle the modality gap by aligning image-to-text pairs, we align image-to-image pairs. To apply such strategies efficiently, parameter-efficient fine-tuning (PEFT) methods like Low-Rank Adaptation (LoRA) (hu2022lora) and prompt tuning (li2021prefixtuning; zhou2022learning) are widely adopted (lester2021the; jahan2025peft). We investigate whether strategies that co-optimize for enhanced representational discriminability alongside generative accuracy enable VLMs to better leverage their internal visual representations.
@@ 3 Experimental setup
Figure 1: An example of the Bongard HOI task. The model must infer a common rule, in this case, "a person is performing a jump on a motorcycle," from the positive examples and determine if the query image follows this rule, which the negative examples do not.
Datasets. We use two Bongard-style datasets: Bongard OpenWorld (wu2024bongardopenworld), split into train (500), validation (100), and test (500) samples with distinct semantic components across splits; and Bongard HOI (jiang2022bongardhoi). For HOI, we use its original splits, with a 4000-sample balanced training set. Its validation ($4\cdot 100$) and test ($4\cdot 200$) sets are categorized by concept novelty (seen/unseen object/action). Each sample contains 6 positive, 6 negative, and 1 query image. For cross-domain generalization, we also evaluate on the Winoground benchmark, a text-image retrieval task designed for testing compositional reasoning.
Models. Baseline performance was established on multiple SoTA VLMs: Phi 3.5 vision 4.2B (abdin2024phi3), Pixtral 12B (agrawal2024pixtral), Gemma3 4B and 27B (team2025gemma), InternVL3 14B (chen2024internvl), and Qwen 2.5 VL 7B and 72B (bai2025qwen25vl). Subsequently, we applied PEFT to Gemma3 4B, Phi and Pixtral.
Evaluation. For all fine-tuning experiments, we saved model checkpoints every epoch and selected the one with the best performance on the validation set. Baselines and best-performing checkpoints were evaluated on their respective test splits. Models demonstrate consistent performance across all conceptual novelty splits (see Appendix L), so we report the average for brevity.
@@ 4 A framework for decomposing VLM reasoning
To diagnose whether VLM failures on abstract reasoning tasks stem from flawed perception or reasoning, we introduce a diagnostic framework. Centered on a non-parametric linear probe, our framework assesses the discriminability of visual embeddings, independent of the generative process, to disentangle representation quality from LLM processing effectiveness. We benchmark performance across various prompt structures, including interleaved vs. labeled image presentation and direct vs. Chain-of-Thought (CoT) prompting (see Appendix C).
@@ 4.1 Methodology
We then extract multi-token image embeddings at two key stages: the vision stage, which captures initial representations from the vision encoder, and the final stage, which captures contextualized representations from the LLM’s final hidden state (detailed in Appendix D). We aggregate these sequences into single vectors, $\vec{v}_{i}$, using mean pooling and L2 normalization. This is a standard technique for producing a single embedding from a sequence of tokens (reimers2019sentencebert), that preserves angular properties relevant to cosine similarity. To quantify the discriminative quality of these embeddings, we probe them using a nearest-centroid classifier. We selected this nearest-centroid strategy as it proved more performant than a nearest-neighbor approach as it mitigates the influence of outlier examples and captures a more robust prototype of the abstract concept (hastie2009the), as shown in Appendix E. We then compute prototype vectors for the positive ($P$) and negative ($N$) example sets by averaging their respective image vectors (e.g., $\vec{c}_{P}=\text{mean}(\vec{v}_{p_{1}},\dots,\vec{v}_{p_{k}})$). A query image ($Q$) is then classified based on its cosine similarity to these centroids (i.e., comparing $\cos(\vec{v}_{Q},\vec{c}_{P})$ to $\cos(\vec{v}_{Q},\vec{c}_{N})$).
4.2 Diagnostic metrics and classification
@@
Our framework dissects VLM performance by statistically comparing three key accuracies: the end-to-end generative accuracy, the linear probe accuracy on initial vision embeddings, and on final contextualized embeddings. Grounded in the linear representation hypothesis, which posits that concepts are encoded as linear structures within a model’s activation space, we define linear probe accuracy on initial vision embeddings as the linear separability ceiling (LSC), quantifying the baseline discriminative power of the visual representations, a primary goal of contrastive pre-training. To classify a model’s performance, we perform a statistical significance test on the difference between its generative accuracy and its LSC, process visualized in Figure 2.
Figure 2: Visual examples of the diagnostic framework and its classifications. The LSC probe (dashed line) is the linear boundary between positive ($\mathbf{c}_{P}$) and negative ($\mathbf{c}_{N}$) centroids. (a) LSC probe: the query ($\mathbf{v}_{Q}$) is correctly classified by the linear probe. (b) Alignment gap: classified when generative accuracy is on average not statistically superior to the LSC ($p\geq 0.05$). The LSC probe succeeds, but the generative model fails. A non-linear model is expected to outperform its linear probe; failure to do so suggests a misalignment. (c) Surpassing the ceiling: classified when generative accuracy is on average statistically superior to the LSC ($p<0.05$). The LSC probe fails, but the generative model succeeds. For illustration, assume reduced dimensions maximally preserve linear properties.
This approach accounts for the uncertainty in both measurements by analyzing the standard error of their difference (burns1981standarderror). A Chi-squared ($\chi^{2}$) test is then used to assess the statistical dependence between paired, trial-by-trial correctness of the model’s generative predictions and the predictions from our linear probes (p < 0.05).
5 Finding: A pervasive alignment gap